# Attosecond X-ray Chronoscopy of Core-level Photoemission

### v3

@@
In conclusion, we have used the novel attosecond capabilities of LCLS to measure core-level photoionization delays of aromatic molecules. These measurements were made possible by the combination of attosecond X-ray pulses with the angular streaking technique. Absolute N-1s photoionization delays were accessed by using the much faster C-1s photoelectrons as a time reference. These self-referenced measurements were realized on a fine energy grid and revealed the presence of local maxima at 3-15 eV above the N-1s threshold. Comparison to advanced core-level photoionization calculations identified two dominant shape resonances in each molecule and showed that photoelectron trapping occurs in the plane of the aromatic molecules. The two dominant mechanisms governing the photoelectron trapping were identified as (i) changes in the barrier height of the molecular potential that lead to an increase of the electron trapping with increasing electronegativity of the functional groups and (ii) the increasing symmetry of the molecular scaffold reduces the coupling between the well-trapped high- and the less-trapped lower- states. Both effects individually and collectively increase the trapping time in the sequence pyridine, pyrimidine, -triazine. These results demonstrate the promise of performing attosecond core-level photoemission experiments at free-electron-laser facilities for the following reasons. The studied molecules are indeed key building blocks of biologically active molecules, as well as molecules employed in optoelectronics, which opens the perspective of studying attosecond electron dynamics in these systems. The key advantage of core-level photoionization delays over valence-shell photoionization delays, is that the photoelectron spectra always remain simple, featuring only one dominant peak per atom of a given type. This will enable the application of the methods described in this work to complex molecules, molecular assemblies and polymers, such as organic semiconductors. Specifically, the effects governing the attosecond electron trapping in aromatic systems unveiled in the present work may become guiding principles for the design and interpretation of experiments studying charge migration and charge transfer in functional materials.
@@ Acknowledgments
Funding: Use of the Linac Coherent Light Source (LCLS), SLAC National Accelerator Laboratory, is supported by the U.S. Department of Energy, Office of Science, Office of Basic Energy Sciences under Contract No. DE-AC02-76SF00515. The work of J.-B.J., X.G., D.H., M.H., S.H. and H.J.W. was supported by ETH Zürich and project 200021_172946 of the Swiss National Science Foundation. The work of M.H. was additionally supported by the European Union’s Horizon 2020 research and innovation program under the Marie Skłodowska-Curie grant agreement No 801459 - FP-RESOMUS. The effort from T.D.D., J.T.O. A.L.W., J.W., T.J.A.W and J.P.C. was supported by DOE, BES, Chemical Sciences, Geosciences, and Biosciences Division (CSGB). Z.G, D.C., J.D., P.L.F., R.R. and A.M. acknowledge support from the Accelerator and Detector Research Program of the Department of Energy, Basic Energy Sciences division. Z.G., P.L.F. and R.R. also acknowledge support from Robert Siemann Fellowship of Stanford University. The effort by R.R.L. and C.W.M. at LBNL was supported by the U.S. DOE BES, CSGB under Contract No. DE-AC02-05CH11231. Calculations at LBNL made use of the resources of the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility, and the Lawrencium computational cluster resource provided by the IT Division at the LBNL. C.S.T. was supported in part by the Berkeley Lab Undergraduate Faculty Fellowship (BLUFF) Program, managed by Workforce Development & Education at Berkeley Lab.
Author Contributions: J.-B.J., T.D., K.U., A.M., J.P.C. and H.J.W. conceived the experiment. All authors contributed to the realization of the experiment and discussed the results. J.-B.J., Z.G. and T.D. developed the data-analysis methodology. J.-B.J. and Z.G. analyzed the data. R.R.L., C.S.T. and C.W.M. performed the calculations of molecular photoionization delays. J.-B.J., Z.G. and H.J.W. wrote the initial draft of the manuscript, which was reviewed and edited by all authors. Competing Interests: The authors declare no competing interests. Data availability statement: All data will be made available on the public ETH data repository upon acceptance of the manuscript.

### v4

@@ V Conclusion
In conclusion, we have used the novel attosecond capabilities of LCLS to measure core-level photoionization delays of aromatic molecules. These measurements were made possible by the combination of attosecond X-ray pulses with the angular streaking technique. Absolute N-1s photoionization delays were accessed by using the much faster and therefore prompt C-1s photoelectrons as a time reference. These self-referenced measurements were realized on a fine energy grid and revealed the presence of local maxima at 3-15 eV above the N-1s threshold. Comparison to advanced core-level photoionization calculations identified two dominant shape resonances in each molecule and showed that photoelectron trapping occurs in the plane of the aromatic molecules. The two dominant mechanisms governing the photoelectron trapping were identified as (i) changes in the barrier height of the molecular potential that lead to an increase of the electron trapping with increasing electronegativity of the functional groups and (ii) the increasing symmetry of the molecular scaffold reduces the coupling between the well-trapped high- and the less-trapped lower- states. Both effects individually and collectively increase the trapping time in -triazine compared to pyridine. These results demonstrate the promise of performing attosecond core-level photoemission experiments at free-electron-laser facilities for the following reasons. The studied molecules are indeed key building blocks of biologically active molecules, as well as molecules employed in optoelectronics, which opens the perspective of extending attosecond chronoscopy to these systems. The key advantage of core-level photoionization delays over valence-shell photoionization delays, is that the photoelectron spectra always remain comparably simple. This will enable the application of the methods described in this work to complex molecules, molecular assemblies and polymers, such as organic semiconductors, which offers the perspective of studying attosecond electron scattering and trapping in functional materials.
VI Methods
Experimental methodology
The attosecond soft X-ray pulses were produced using the photocathode shaping method [40]. The mean pulse energy is J and the mean FWHM bandwidth is eV. From our previous measurements, we can estimate a FWHM temporal pulse duration of as for the X-ray pulses.
The attosecond soft X-ray pulses described above were focused to a spot-size of m in diameter using a pair of Kirkpatrick-Baez focusing mirrors [41]. The -nm streaking pulse with a duration of 40 fs was produced by a commercial optical parametric amplifier (TOPAS-HE, Light Conversion) which is pumped by a fs, mJ, -nm laser source. The signal beam was sent through a superachromatic quarter-wave plate (Thorlabs SAQWP05M-1700) to produce a circularly polarized laser field and was focused with a mm CaF2 lens before being reflected from a silver-coated mirror with a -mm-diameter hole to co-propagate the laser and X-ray fields to the interaction point in the c-VMI spectrometer chamber.
High pressure (6.4 bar) helium gas was bubbled through a vessel heated to 80-140∘C to deliver the samples into the c-VMI spectrometer chamber, and the photoelectrons ionized from the samples were collected using an electrostatic lens. A variable-line spaced (VLS) grating X-ray photon spectrometer [42, 43] was employed downstream of the interaction point to obtain the spectra for individual X-ray shots.
Data analysis
The relative photoionization delay between the carbon and nitrogen K-shell photoemission features, , is related to the difference in streaking angle, , between the momentum shift of the nitrogen and carbon photoemission features, , where fs is the period of the IR streaking laser. We use two methods to extract the differential streaking angle from the dataset. The first method uses a partial covariance approach to isolate the correlated changes of streaking-induced variations in the photoelectron momentum distribution, while the second procedure uses a sorting method which tags the arrival time of the attosecond X-ray pulse relative to the IR-field. The two approaches are further illustrated and compared in the supplementary information.
Theoretical methodology
The N photoionization of the three azabenzenes considered here were studied using two methods, the basis-set complex Kohn method [44] and ePolyScat [45, 46]. The one-electron basis set used to obtain the occupied molecular orbitals was a correlation consistent valence triple zeta [47] and the geometry used minimized the Hartree-Fock energy. The ePolyScat calculations used a static-exchange potential with equivalent-core method. An additional CP potential was introduced to approximate the interaction between the emitted electron and the residual ion of the molecule [48]. The calculated values with and without the CP potential are compared in the SM. The multichannel calculation from the coupled N holes using the complex-Kohn method [49] and the one-channel calculation using ePolyScat from the localized N hole show consistency, as demonstrated in the SM.
Data availability
All data will be made available on the public ETH data repository upon acceptance of the manuscript.
@@ Acknowledgments
Use of the Linac Coherent Light Source (LCLS), SLAC National Accelerator Laboratory, is supported by the U.S. Department of Energy, Office of Science, Office of Basic Energy Sciences under Contract No. DE-AC02-76SF00515. The work of J.-B.J., X.G., D.H., M.H. (ETH Zürich), S.H. and H.J.W. was supported by ETH Zürich and project 200021_172946 of the Swiss National Science Foundation. The work of M.H. (ETH Zürich) was additionally supported by the European Union’s Horizon 2020 research and innovation program under the Marie Skłodowska-Curie grant agreement No 801459 - FP-RESOMUS. The effort from T.D., J.T.O. A.L.W., J.W., T.J.A.W. and J.P.C. was supported by DOE, BES, Chemical Sciences, Geosciences, and Biosciences Division (CSGB). Z.G., D.C., J.D., P.L.F., R.R.R. and A.M. acknowledge support from the Accelerator and Detector Research Program of the Department of Energy, Basic Energy Sciences division. Z.G., P.L.F. and R.R.R. also acknowledge support from Robert Siemann Fellowship of Stanford University. The effort by R.R.L. and C.W.M. at LBNL was supported by the U.S. DOE BES, CSGB under Contract No. DE-AC02-05CH11231. Calculations at LBNL made use of the resources of the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility, and the Lawrencium computational cluster resource provided by the IT Division at the LBNL. C.S.T. was supported in part by the Berkeley Lab Undergraduate Faculty Fellowship (BLUFF) Program, managed by Workforce Development & Education at Berkeley Lab.
Author contributions
J.-B.J., T.D., K.U., A.M., J.P.C. and H.J.W. conceived the experiment. All authors contributed to the realization of the experiment and discussed the results. J.-B.J., Z.G. and T.D. developed the data-analysis methodology. J.-B.J. and Z.G. analyzed the data. R.R.L., C.S.T. and C.W.M. performed the calculations of molecular photoionization delays. J.-B.J., Z.G. and H.J.W. wrote the initial draft of the manuscript, which was reviewed and edited by all authors.
Competing interests
The authors declare no competing interests.